# MonoMAE: Enhancing Monocular 3D Detection through Depth-Aware Masked Autoencoders

### v1

@@ 4.2 Benchmarking with the State-of-the-Art
Quantitative Benchmarking. We conducted quantitative benchmarking over the Car category of the test set on KITTI 3D as shown in Table 1, where all evaluations were performed on the official online test server [13] for fairness. We can see that MonoMAE achieves superior detection performance consistently across all metrics, without using any extra training data such as image depths, video sequences, LiDAR points, and CAD 3D models. In addition, MonoMAE outperforms more for the Moderate and Hard categories where various occlusions happen much more frequently than the Easy category. The superior performance is largely attributed to our designed depth-aware masking and completion network which masks queries to simulate object occlusions in the feature level and reconstructs the masked queries to learn occlusion-tolerant visual representations, respectively.
@@
Table 2: Ablation study of technical designs in MonoMAE on the KITTI 3D val set. ‘NOQG’, ‘DAM’, and ‘CN’ denote the Non-Occluded Query Grouping, the Depth-Aware Masking, and the Completion Network, respectively. The symbol * indicates the baseline. The best results are highlighted in bold.
@@
Table 3: Ablation study of three masking strategies on the KITTI 3D val set. The best results are highlighted in bold.
@@
Qualitative Benchmarking. Figure 5 shows qualitative benchmarking on the KITTI 3D val set. It can be observed that compared with two state-of-the-art methods GUPNet and MonoDETR, our MonoMAE produces more accurate 3D detection consistently for both non-occluded and occluded objects, even for some hard scenarios like distant objects. Specifically, GUPNet and MonoDETR tend to miss the detection of highly occluded object in Cases 1 and 2 as highlighted by red arrows, while our MonoMAE performs clearly better by detecting those challenging objects successfully, demonstrating its superior capability on handling object occlusions. More visualization results can be found in supplementary materials.
@@ 4.3 Ablation Study
We conduct extensive ablation studies to examine our proposed MonoMAE. Specifically, we examine MonoMAE from the aspect of the technical designs, query mask approaches, as well as loss functions.
Network Designs. We examine the effectiveness of two key designs in MonoMAE, namely, the Depth-Aware Masking module (DAM) and the Completion Network (CN) (on the test set of KITTI 3D). We formulate the baseline by including the Non-Occluded Query Grouping module (NOQG), which does not affect the network training as both identified occluded and non-occluded queries are fed to train 3D detectors. While incorporating DAM only in Rows 2 and 4, the 3D detection degrades clearly as queries are masked but not reconstructed which leads to further information loss. While incorporating CN alone in Rows 3 and 5, the detection improves clearly as the completion helps learn better representations for naturally occluded queries. In addition, incorporating DAM and CN on top of NOQG in Row 7 performs clearly better than incorporating DAM and CN alone in Row 6, as the former applies masking and completion to non-occluded queries only. It also shows that masking naturally occluded queries to train the completion network is harmful to the learned representations.
Masking Strategies. We examine how different masking strategies affect monocular 3D detection. We studied three masking strategies as shown in Table 3. The first strategy masks the input images randomly, aiming to assess the value of masking and completing at the feature instead of image level. We can observe that the image-level masking yields clearly lower performance as compared with query masking in the feature space, largely due to the complication in masking and reconstructing images with a lightweight completion network. The second strategy masks query features randomly without considering object depths, aiming to evaluate the importance of object depths in query masking. The experiments show that random query masking outperforms the image-level masking significantly. The third strategy performs the proposed depth-aware query masking. It outperforms random query masking consistently, demonstrating the value of incorporating object depth information in query masking.
Table 4: Ablation study of the loss functions on the KITTI 3D val set. $L_{occ}$ and $L_{com}$ refer to the occlusion classification loss and the completion loss, respectively. The best results are highlighted in bold.
@@
Loss Functions. We examine the impact of the occlusion classification loss $L_{occ}$ and the completion loss $L_{com}$ in Equation 2 and 6, where $L_{occ}$ supervises the occlusion classification network (in Non-Occluded Query Grouping) to predict whether the queries are occluded and $L_{com}$ supervises the Completion Network to reconstruct the masked queries. As Table 4 shows, while implementing $L_{occ}$ alone, the occlusion prediction is supervised while the query reconstruction is unsupervised. The network under such an objective does not learn well as the Completion Network cannot reconstruct object queries well without sufficient supervision. While implementing $L_{com}$ alone, the occlusion classification network cannot identify occluded and non-occluded queries accurately where many occluded queries are fed for masking, leading to more query occlusion and poor detection performance. While employing both losses concurrently, the performance improves significantly as non-occluded queries can be identified for masking and reconstruction, leading to occlusion-tolerant representations.
@@ 4.4 Discussions
Efficiency Comparison. We compare the inference time of several representative monocular 3D detection methods on the KITTI val set, where all compared methods are evaluated with one NVIDIA V100 GPU under the same computational environment for fairness. As Table 5 shows, GUPNet, MonoDTR, and MonoDETR have an average inference time of 40ms, 37ms, and 43ms for each image, respectively. As a comparison, the proposed MonoMAE takes the shortest inference time, demonstrating its good efficiency in monocular 3D detection. The good efficiency can be attributed to several factors. First, MonoMAE does not involve Depth-Aware Masking but just the Completion Network for reconstructing occluded queries. In addition, the Completion Network is lightweight which just introduces 2ms computational overhead as shown in the last two columns of Table 5. Further, we analyzed the Completion Network in terms of network parameters and floating-point operations per second (FLOPs), showing that it has very limited 2.22G parameters and 0.08M in FLOPs.
Generalization Ability. We examine the generalization capability of the proposed MonoMAE by directly applying the KITTI-trained MonoMAE model to the car Category of the nuScenes validation set without additional training. The performance on the KITTI validation set is also reported for reference. Table 6 illustrates that MonoMAE attains the highest or second-highest performance across various metrics on the nuScenes frontal validation set. This indicates that despite the domain shift from KITTI to nuScenes, MonoMAE still maintains satisfactory performance. DEVIANT [22] also achieves great monocular 3D detection performance on nuScenes as it treats objects of different depths equally and is thus robust to depth variations.
Figure 6: Completion Network validation with loss visualization. The red curve shows the loss of the original queries and completed queries by the Completion Network. The blue curve shows the loss of the original queries and the masked queries without completion. Completing the masked queries by the proposed Completion Network improves the learning significantly.
Effect of Completion. We also validate the effect of completion by comparing the loss of the completed queries and the original queries (red) with the loss of the masked queries and the original queries (blue). As Figure 6 shows, the training loss while including the Completion Network drops much faster than the training loss without including the Completion Network, demonstrating that the Completion Network helps acquire occlusion-tolerant representations effectively by learning to reconstruct the masked queries.
5 Limitations
The proposed MonoMAE could be improved in several aspects. Specifically, MonoMAE leverages depth-aware masking to mask non-occluded queries to simulate object occlusions in the feature space. However, the masked queries may have different patterns as compared with the features of naturally occluded object queries. Such a gap could affect the mask reconstruction and the completed queries which further affects the learned representations and monocular 3D detection performance. This issue could be mitigated by introducing generative networks that learn distributions from extensive real-world data for generating occlusion patterns that are more similar to natural occlusions. In addition, MonoMAE could be evaluated and analyzed more extensively over larger datasets when they are available.
@@ 6 Conclusion
This paper presents MonoMAE, a novel method inspired by the Masked Autoencoders (MAE) to deal with the pervasive occlusion problem in the monocular 3D object detection task. MonoMAE consists of two key designs. The first is a depth-aware masking module, which simulates the occlusion for non-occluded object queries in the feature level during training. The second is a lightweight completion network, which reconstructs and completes the masked object queries. Quantitative and qualitative experiment results show that MonoMAE learns enhanced 3D representations and achieves superior monocular 3D detection performance for both occluded and non-occluded objects. Moving forward, we plan to investigate generative approaches to simulate natural occlusion patterns for various 3D detection tasks.

### v2

@@ 4.2 Benchmarking with the State-of-the-Art
Quantitative Benchmarking. Table 1 shows quantitative experiments on the test set of dataset KITTI 3D, where all evaluations were performed on the official online test server [13] for fairness. We can see that MonoMAE achieves superior detection performance consistently across all metrics, without using any extra training data such as image depths, video sequences, LiDAR points, and CAD 3D models. In addition, MonoMAE outperforms more for the Moderate and Hard categories where various occlusions happen much more frequently than the Easy category. The superior performance is largely attributed to our designed depth-aware masking and completion network, which masks queries to simulate object occlusions in the feature space and reconstructs the masked queries to learn occlusion-tolerant visual representations, respectively.
@@
Table 2: Ablation study of technical designs in MonoMAE on the KITTI 3D val set. ‘NOQG’, ‘DAM’, and ‘CN’ denote Non-Occluded Query Grouping, Depth-Aware Masking, and Completion Network, respectively. The symbol * indicates the baseline. The best results are highlighted in bold.
@@
Table 3: Ablation study of masking strategies on the KITTI 3D val set. The best results are in bold.
@@
Qualitative Benchmarking. Figure 4 shows qualitative benchmarking on the KITTI 3D val set. It can be observed that compared with two state-of-the-art methods GUPNet and MonoDETR, the proposed MonoMAE produces more accurate 3D detection consistently for both non-occluded and occluded objects, even for challenging scenarios like distant objects. Specifically, GUPNet and MonoDETR tend to miss the detection of highly occluded object in Cases 1 and 2 as highlighted by red arrows. As a comparison, MonoMAE performs clearly better by detecting those challenging objects successfully, demonstrating its superior capability on handling object occlusions.
@@ 4.3 Ablation Study
We conduct extensive ablation studies to examine the proposed MonoMAE. Specifically, we examine MonoMAE from the aspect of the technical designs, query masking strategies, as well as loss functions.
Network Designs. We examine the effectiveness of two key designs in MonoMAE, namely, the Depth-Aware Masking module (DAM) and the Completion Network (CN) (on the validation set of KITTI 3D), as shown in Table 2. We formulate the baseline by including the Non-Occluded Query Grouping module (NOQG), which does not affect the network training as both identified occluded and non-occluded queries are fed to train 3D detectors. When CN is not used in Rows 2 and 4, the 3D detection degrades as queries are masked but not reconstructed which leads to further information loss. While not incorporating DAM in Rows 3 and 5, the detection improves clearly compared with the baseline, as the completion helps learn better representations for naturally occluded queries. In addition, incorporating DAM and CN on top of NOQG in Row 7 performs clearly better than incorporating DAM and CN alone in Row 6, as the former applies masking and completion to non-occluded queries only. It also shows that masking naturally occluded queries to train the completion network is harmful to the learned representations.
Masking Strategies. We examine how different masking strategies affect monocular 3D detection. We studied three masking strategies as shown in Table 3. The first strategy masks the input images randomly, aiming to assess the value of masking and completing in the feature instead of image space. We can observe that the image-level masking yields clearly lower performance as compared with query masking in the feature space, largely due to the complication in masking and reconstructing images with a lightweight completion network. The second strategy masks query features randomly without considering object depths, aiming to evaluate the importance of object depths in query masking. The experiments show that random query masking outperforms the image-level masking significantly. The third strategy performs the proposed depth-aware query masking. It outperforms the feature-space random masking consistently, demonstrating the value of object depths for query masking.
Table 4: Ablation study of the loss functions on the KITTI 3D val set. $L_{occ}$ and $L_{com}$ refer to the occlusion classification loss and the completion loss, respectively. The best results are in bold.
@@
Loss Functions. We examine the impact of the occlusion classification loss $L_{occ}$ and the completion loss $L_{com}$ in Equations 2 and 6, where $L_{occ}$ supervises the occlusion classification network (in Non-Occluded Query Grouping) to predict whether the queries are occluded and $L_{com}$ supervises the Completion Network to reconstruct the masked queries. As Table 4 shows, while implementing $L_{occ}$ alone, the occlusion prediction is supervised while the query reconstruction is unsupervised. The network under such an objective does not learn well as the Completion Network cannot reconstruct object queries well without sufficient supervision. While implementing $L_{com}$ alone, the occlusion classification network cannot identify occluded and non-occluded queries accurately where many occluded queries are fed for masking, leading to more query occlusion and poor detection performance. While employing both losses concurrently, the performance improves significantly as non-occluded queries can be identified for masking and reconstruction, leading to occlusion-tolerant representations.
@@ 4.4 Discussions
Efficiency Comparison. We compare the inference time of several representative monocular 3D detection methods on the KITTI val set, where all compared methods are evaluated with one NVIDIA V100 GPU under the same computational environment for fairness. As Table 5 shows, GUPNet, MonoDTR, and MonoDETR have an average inference time of 40ms, 37ms, and 43ms for each image, respectively. As a comparison, the proposed MonoMAE takes the shortest inference time, demonstrating its good efficiency in monocular 3D detection. Further, we analyzed the Completion Network in terms of network parameters and floating-point operations per second (FLOPs), showing it has very limited 2.22G parameters and 0.08M in FLOPs.
Generalization Ability. We examine the generalization capability of the proposed MonoMAE by directly applying the KITTI-trained MonoMAE model to the car Category of the nuScenes validation set without additional training. The detection performance on the KITTI validation set is also reported for reference. Table 6 shows that MonoMAE attains the highest or second-highest detection performance across various metrics on the nuScenes frontal validation set. This indicates that despite the domain shift from KITTI to nuScenes, MonoMAE still maintains satisfactory performance. Since DEVIANT [24] is equivariant to the depth translations, it sometimes has higher performance.
@@ 5 Conclusion
This paper presents MonoMAE, a novel method inspired by the Masked Autoencoders (MAE) to deal with the pervasive occlusion problem in the monocular 3D object detection task. MonoMAE consists of two key designs. The first is a depth-aware masking module, which simulates the occlusion for non-occluded object queries in the feature space during training. The second is a lightweight completion network, which reconstructs and completes the masked object queries. Quantitative and qualitative experiment results show that MonoMAE learns enhanced 3D representations and achieves superior monocular 3D detection performance for both occluded and non-occluded objects. Moving forward, we plan to investigate generative approaches to simulate natural occlusion patterns for various 3D detection tasks.
Limitations. MonoMAE leverages depth-aware masking to mask non-occluded queries to simulate object occlusions in the feature space. However, the masked queries may have different patterns as compared with the features of naturally occluded object queries. Such a gap could affect the reconstruction of complete queries and monocular 3D detection performance. This issue could be mitigated by introducing generative networks that learn distributions from extensive real-world data for generating occlusion patterns that are more similar to natural occlusions.
Acknowledgments and Disclosure of Funding
This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).